# Multi-Agent Taskforce Collaboration: Self-Correction of Compounding Errors in Long-Form Literature Review Generation

### v1

@@ Abstract
Literature reviews play an important role in scientific research. Recent advances in large language models (LLMs) have boosted the development of automated systems for the entire literature review workflow, from retrieval to manuscript drafting. However, a key challenge is that mistakes made in early stages can propagate and amplify in subsequent steps, leading to compounding errors that undermine the faithfulness of the final review. To tackle this issue, we propose the Multi-Agent Taskforce Collaboration (MATC) framework, which consists of a manager agent and four executor agents for literature searching, outline generation, fact localization, and manuscript drafting. We propose three novel collaboration paradigms, forming exploration, exploitation, and experience taskforces, to effectively organize agents and mitigate compounding errors both between and within executor agents. Experimental results show that MATC achieves state-of-the-art performance on existing benchmarks. We further propose a new benchmark dataset featuring more diverse topics for faithful literature review generation.
Introduction
@@
Literature reviews play an important role in scientific research by synthesizing existing knowledge to identify research gaps, establish theoretical frameworks, and guide future studies (Ermel et al. 2021; Liu et al. 2023). Over recent decades, the rapid growth in the number of scientific publications has made it increasingly challenging for researchers to efficiently retrieve, organize, and synthesize relevant information in their fields (Wang et al. 2024b). In response, automatic literature review systems that leverage advances in artificial intelligence have become increasingly valuable, as they help reduce manual workload, provide timely coverage of new research, and facilitate the identification of key trends and research gaps within the extensive literature (Zhang et al. 2025b; Wang et al. 2024b; Liu et al. 2023).
@@
Early approaches primarily leverage multi-document summarization techniques to generate summaries from a set of reference papers (Hoang and Kan 2010; Hu and Wan 2014; Erera et al. 2019), enabling the automatic creation of related work sections given a set of references as input. In recent years, large language models (LLMs) demonstrate impressive planning and reasoning abilities (Wang et al. 2024a; Zhao et al. 2024; Zhong and Wang 2024; Nandy and Bandyopadhyay 2025; Zhang et al. 2025b). This leads to a growing research area that employs LLMs as controllers to construct autonomous agents capable of effectively performing complex tasks (Wang et al. 2024a). Building on these advances, recent state-of-the-art research, e.g., AutoSurvey (Wang et al. 2024b) and SurveyX (Liang et al. 2025), achieves remarkable success by designing LLM-based autonomous agents to automate the entire pipeline of literature review and generate long-form manuscripts.
@@
Although these advances enable the automatic generation of human-like literature reviews with reduced manual effort, previous studies on multi-step models show that the inherent long-horizon workflow of such systems inevitably introduces the risk of compounding errors. In the workflow, small mistakes made at earlier steps can accumulate and propagate through subsequent steps, as errors that occur at a certain step will affect the next state (Wang et al. 2025a). Without effective self-correction mechanisms, these deviations will further expand the error in later steps (Han and Zhang 2023; Venkatraman, Hebert, and Bagnell 2015; Hejna, Abbeel, and Pinto 2023).
@@
This paper investigates the phenomenon of error accumulation in automatic literature review generation. As a demonstration, we employ the state-of-the-art large language model GPT-4.1 (OpenAI 2025) to conduct a literature review generation workflow on the topic of EEG (electroencephalogram)-based emotion recognition, with a target length of 4,000 words. As shown in Fig. 1, when the retrieval stage fails to identify relevant literature, the resulting errors can mislead the outlining process. An irrelevant outline may further cause the drafted manuscript to deviate from the intended topic. In long-form text generation, the increased complexity and the introduction of more dependent steps further exacerbate the compounding errors.
@@
To tackle this challenge, we propose the Multi-Agent Taskforce Collaboration (MATC) framework. MATC consists of two types of autonomous agents: the manager agent and the executor agents. The manager agent orchestrates the workflow, while the executor agents each specialize in a distinct step of the process, such as literature searching, outline generation, fact localization, or manuscript drafting. Inspired by practices in organizational quality management (Hackman and Wageman 1995), the MATC framework introduces the first multi-agent collaboration algorithm that orchestrates agents organized as taskforces. A taskforce refers to a temporary group of specialized agents focused on specific subtasks (Altier 1987; Aikin 1949; de Waard and Kramer 2008).
Given a user’s instruction, a manager agent oversees the entire workflow and initiates the formation of different taskforces. To explore the research field, the senior manager organizes an exploration taskforce. Since compounding errors may accumulate during exploration between the searching and outlining agents, we design a self-correcting tree that begins with a broad, low-barrier overview and incrementally refines both the literature selection and outline in a hierarchical manner.
To exploit the retrieved literature for manuscript generation, the locating and drafting agents form an exploitation taskforce. As compounding errors may also accumulate during exploitation between the locating and drafting agents, we design a self-correcting cycle. In this cycle, draft sections and extracted fact snippets are iteratively refined through alternating optimization, effectively bridging the draft to the underlying literature.
To mitigate compounding errors that may occur within each functional step due to unpredictable agent-environment interactions, the senior manager agent creates experience taskforces for each agent type. These taskforces maintain the historical experience of agents and implement a dynamic checklist to guide self-correction and revision. Finally, the senior manager integrates the outputs of all taskforces to produce the final manuscript.
Related Work
@@
Automating the summarization of research papers has long attracted interest. Early work mainly used multi-document summarization techniques. In 2010, Hoang et al. introduced automatic related work summarization, extracting sentences from multiple articles to form an extractive summary tailored to the target paper (Hoang and Kan 2010). Lu et al. constructed Multi-XScience, a large-scale dataset for generating related work sections based on a paper’s abstract and cited articles, serving as a benchmark for later studies (Lu, Dong, and Charlin 2020). Chen et al. proposed a relation-aware multi-document encoder and a relation graph to capture cross-document dependencies for abstractive related work generation (Chen et al. 2021). Xiao et al. developed PRIMERA, a pre-trained encoder-decoder model with novel objectives for linking and aggregating information across documents (Xiao et al. 2022). Liu et al. proposed aligning sentences by category and using sparse transformers to organize information from many references (Liu et al. 2023). Zhang et al. identified relations between concepts from literature and automatically built knowledge minigraphs to organize concepts and generate logical multi-document summaries (Zhang et al. 2025b). These methods advance paragraph-level generation given reference sets. However, fully automated literature review paper generation requires further abilities, such as outline planning and reference retrieval.
@@
Recently, advances in LLMs have greatly improved the scope and quality of automated literature reviews, enabling end-to-end systems to generate literature review papers. Wang et al. introduced AutoSurvey, a framework with retrieval, outline generation, parallel subsection drafting, integration, and rigorous evaluation (Wang et al. 2024b). Shao et al. developed STORM, which models the pre-writing stage by discovering diverse perspectives, simulating expert conversations, and curating information to create an outline used to generate a full Wikipedia article (Shao et al. 2024). Liang et al. designed SurveyX, an end-to-end solution for automated survey generation, covering online literature search, organization, and survey writing (Liang et al. 2025). While these methods represent impressive progress, little research has addressed the risk of errors in generated outputs. Self-correction algorithms in LLMs are not suitable for the workflow of literature review generation (Madaan et al. 2023; Shinn et al. 2023; Ferraz et al. 2024). As systems become more complex, error accumulation becomes a significant challenge.
Methodology
@@
In detail, given a user instruction $U$, the senior manager $A_{M}$ initiates the exploration taskforce. The senior manager constructs a tree with the user instruction as the root node, representing the overall direction for exploring the field. This tree is expanded by identifying sub-directions of interest. For the root node, $A_{M}$ first invokes the searching agent to retrieve literature based on the user instruction, resulting in a set of relevant papers. The searching agent $A_{S}$ operates as follows:
@@
Next, the senior manager assigns the outlining agent to determine sub-directions based on the titles and abstracts of the retrieved literature. The outlining agent $A_{O}$ operates as follows:
@@
For each outline node $O_{j}^{(d)}$ at depth $d$, the senior manager decides whether further decomposition is needed according to the following criterion:
@@
In detail, given the outline tree $\mathcal{T}$ produced as mentioned above, the senior manager $A_{M}$ initiates the exploitation taskforce. For each leaf node in the tree, $A_{M}$ identifies its parent outline node, denoted as $\{\mathcal{O}_{j}\}$, and the associated literature set $\{\mathcal{L}_{i}\}$. The exploitation taskforce then proceeds through iterative cycles. At the initial iteration $t=0$, the locating agent $A_{L}$ extracts relevant information based on the outline node $\{\mathcal{O}_{j}\}$ and the literature set $\{\mathcal{L}_{i}\}$, and operates as follows:
@@
where $\psi(\cdot,\cdot)$ denotes the similarity between consecutive iterations of the draft, measured by the ROUGE-1 score (Lin 2004), $\epsilon$ is a similarity threshold that indicates no further significant improvements can be made, and $t_{\max}$ is the maximum number of allowed iterations.
The final content $M^{(t^{*})}_{i}$, where $t^{*}$ denotes the stopping iteration for parent node $i$, serves as the final draft. Finally, the senior manager $A_{M}$ aggregates the finalized drafts for all sections and delivers the completed literature review to the user.
Experience Taskforce
The experience taskforce is designed to execute tasks assigned by the senior manager, taking into account the risk of deviations at each step arising from unpredictable agent-environment interactions. To mitigate compounding errors, we propose an experience-based strategy.
In detail, for each agent $A_{W}$, the senior manager agent $A_{M}$ initiates the experience taskforce. The agent first generates the primary results, then $A_{M}$ leverages historical experience to review these results, after which $A_{W}$ refines its outputs based on the feedback provided. At the initial step, $A_{W}$ produces the first output solely based on the context:
@@
where $Y^{(0)}$ is the initial output, and $C$ is the context fetched by the senior manager agent.
Survey Length Methods Citation Quality Content Quality Recall $\uparrow$ Precision $\uparrow$ Coverage $\uparrow$ Structure $\uparrow$ Relevance $\uparrow$ Avg. $\uparrow$ 8k Naive RAG (2024) 78.14 71.92 4.40 3.86 4.86 4.37 AutoSurvey (2024) 82.48 77.42 4.60 4.46 4.80 4.62 SurveyX (2025) - - - - - - Proposed 98.17 89.28 4.97 4.95 5.00 4.97 16k Naive RAG (2024) 71.48 65.31 4.46 3.66 4.73 4.28 AutoSurvey (2024) 81.34 76.94 4.66 4.33 4.86 4.62 SurveyX (2025) - - - - - - Proposed 98.11 88.52 4.98 4.95 4.98 4.97 32k Naive RAG (2024) 79.88 65.03 4.41 3.75 4.66 4.27 AutoSurvey (2024) 83.14 78.04 4.73 4.26 4.80 4.60 SurveyX (2025) - - - - - - Proposed 98.03 87.92 4.98 4.93 4.98 4.96 64k Naive RAG (2024) 68.79 61.97 4.40 3.66 4.66 4.24 AutoSurvey (2024) 82.25 77.41 4.73 4.33 4.86 4.64 SurveyX (2025) 85.14 78.02 4.95 4.90 4.95 4.93 Proposed 97.27 87.90 5.00 4.91 4.98 4.96
@@
The senior manager agent then recalls historical experiences $\mathcal{E}$. At each iteration $x$, the senior manager agent selects the most useful review from $\mathcal{E}$:
@@
Here, the senior manager agent records the number of subsequent revision comments for each review. The most useful review is defined as the one that has led to the fewest further comments, where $\mathcal{R}_{\min}=\{r\in\mathcal{R}:\text{revision}(r)=r_{\min}\}$, $\text{revision}(r)$ denotes the number of revisions after using review $r$, and $\text{usage}(r)$ represents the usage count of review $r$.
The senior manager agent then provides comments on the worker’s output:
@@
where $R^{(x)}$ denotes the comments at iteration $x$.
For subsequent iterations ($x\geq 1$), the worker agent refines the output with the input and the latest comments:
@@
The experience taskforce continues until the output meets the senior manager’s requirements and no further comments are provided, or until a maximum number of iterations is reached. In the following, we detail how workers cooperate with the senior manager.
The searching agent formulates search queries based on user instructions, submits queries to search engines, and collects relevant literature. It filters and ranks the retrieved references by citation count, removes incomplete entries, and selects the top $Q_{k}$ unique references for further use. The outlining agent utilizes the gathered literature and user instructions to prompt an LLM for generating sub-outlines. The locating agent downloads and parses full-text papers, applies LLM-based previewing to select pertinent pages, and extracts concise text snippets for drafting. The drafting agent prompts the LLM to synthesize coherent text, by incorporating selected references and prior drafts.
Experiments
Experimental Settings
We conduct experiments on two benchmark datasets, AutoSurvey (Wang et al. 2024b) and SurveyEval (Wang et al. 2025b), as well as a self-constructed benchmark dataset, TopSurvey. For hyperparameters, we set the maximum number of iterations for exploration, exploitation, and experience taskforces to 4. The hyperparameter $\theta$ is set to 500. We use GPT-4.1 as the LLM for both generation and evaluation. The evaluation comprises two categories of metrics. For citation quality, we adopt citation recall and citation precision as proposed by Wang et al. (2024b): recall measures whether cited passages fully support all statements, while precision measures the proportion of relevant citations that support their corresponding statements. For content quality, following Wang et al. (2024b), we use coverage, structure, and relevance, each rated by LLMs on a 5-point scale. Coverage assesses the extent to which the survey addresses all relevant aspects of the topic; structure evaluates logical organization and coherence; and relevance measures alignment with the specified research topic. We do not filter out fractional scores, such as 4.5.
Comparision Experiments

### v2

@@ Abstract
Compounding error is critical in long-form literature review generation, where minor inaccuracies cascade and amplify across subsequent steps, severely compromising the faithfulness of the final output. To address this challenge, we propose the Multi-Agent Taskforce Collaboration (MATC) framework, which proactively mitigates errors by orchestrating LLM-based agents into three specialized taskforces: (1) an exploration taskforce that interleaves retrieval and outlining using a tree-based strategy to establish a grounded structure; (2) an exploitation taskforce that iteratively cycles between fact location and draft refinement to ensure evidential support; and (3) a feedback taskforce that leverages historical experience for self-correction before errors propagate. Experimental results show that MATC achieves state-of-the-art performance on existing benchmarks (AutoSurvey and SurveyEval), significantly outperforming strong baselines in both citation quality (e.g., +15.7% recall) and content quality. We further contribute TopSurvey, a new large-scale benchmark of 195 peer-reviewed survey topics, on which MATC maintains robust performance, demonstrating its generalizability.
Multi-Agent Taskforce Collaboration: Self-Correction of Compounding Errors in Long-Form Literature Review Generation
Zhi Zhang1, Yan Liu1thanks: Corresponding author., Zhejing Hu1, Gong Chen2, Sheng-hua Zhong3, Jiannong Cao1 1Department of Computing, The Hong Kong Polytechnic University, Hong Kong, China 2FireTorch Partners, Shenzhen, China 3College of Computer Science and Software Engineering, Shenzhen University, Shenzhen, China {zhi271.zhang, zhejing.hu}@connect.polyu.hk, {yan.liu, jiannong.cao}@polyu.edu.hk, heinz@clozzz.com, csshzhong@szu.edu.cn
1 Introduction
@@
Literature reviews play an important role in scientific research by synthesizing existing knowledge to identify research gaps, establish theoretical frameworks, and guide future studies Ermel et al. (2021); Liu et al. (2023). Over recent decades, the rapid growth in the number of scientific publications has made it increasingly challenging for researchers to efficiently retrieve, organize, and synthesize relevant information in their fields Wang et al. (2024b). In response, automatic literature review systems that leverage advances in artificial intelligence have become increasingly valuable, as they help reduce manual workload, provide timely coverage of new research, and facilitate the identification of key trends and research gaps within the extensive literature Zhang et al. (2025); Wang et al. (2024b); Liu et al. (2023).
@@
Early approaches primarily leverage multi-document summarization techniques to generate summaries from a set of reference papers Hoang and Kan (2010); Hu and Wan (2014); Erera et al. (2019), enabling the automatic creation of related work sections given a set of references as input. In recent years, large language models (LLMs) demonstrate impressive planning and reasoning abilities Wang et al. (2024a); Zhao et al. (2024); Zhong and Wang (2024); Nandy and Bandyopadhyay (2025); Zhang et al. (2025). This leads to a growing research area that employs LLMs as controllers to construct autonomous agents capable of effectively performing complex tasks Wang et al. (2024a). Building on these advances, recent state-of-the-art research, e.g., AutoSurvey Wang et al. (2024b) and SurveyX Liang et al. (2025), achieves remarkable success by designing LLM-based autonomous agents to automate the entire pipeline of literature review and generate long-form manuscripts.
@@
Although these advances enable the automatic generation of human-like literature reviews with reduced manual effort, previous studies on multi-step models show that the inherent long-horizon workflow of such systems inevitably introduces the risk of compounding errors. In the workflow, small mistakes made at earlier steps can accumulate and propagate through subsequent steps, as errors that occur at a certain step will affect the next state Wang et al. (2025a). Without effective self-correction mechanisms, these deviations will further expand the error in later steps Han and Zhang (2023); Venkatraman et al. (2015); Hejna et al. (2023).
@@
This paper investigates the phenomenon of error accumulation in automatic literature review generation. As a demonstration, we employ the state-of-the-art large language model GPT-4.1 OpenAI (2025) to conduct a literature review generation workflow on the topic of EEG (electroencephalogram)-based emotion recognition, with a target length of 4,000 words. As shown in Fig. 1, when the retrieval stage fails to identify relevant literature, the resulting errors can mislead the outlining process. An irrelevant outline may further cause the drafted manuscript to deviate from the intended topic. In long-form text generation, the increased complexity and the introduction of more dependent steps further exacerbate the compounding errors.
@@
To tackle this challenge, we propose the Multi-Agent Taskforce Collaboration (MATC) framework. MATC consists of two types of autonomous agents: the manager agent and the executor agents. The manager agent orchestrates the workflow, while the executor agents each specialize in a distinct step of the process, such as literature searching, outline generation, fact localization, or manuscript drafting.
Inspired by organizational quality management Hackman and Wageman (1995), we propose to orchestrate agents organized as taskforces. A taskforce refers to a temporary group of specialized agents focused on specific subtasks Altier (1987); Aikin (1949); de Waard and Kramer (2008). We formulate three types of taskforces, each designed with specific error mitigation mechanisms.
Given a user’s instruction, the manager agent oversees the entire workflow and initiates the formation of different taskforces. To mitigate error accumulation between searching and outlining agents, the manager organizes an exploration taskforce employing a tree-based strategy. The taskforce begins with a broad overview and incrementally determines the literature and outline at each level, preventing errors in ungrounded outlines or biased retrieval. To mitigate error accumulation between locating and drafting agents, the manager organizes an exploitation taskforce employing an iterative cycle: the draft guides fact location from the literature, while the located facts inform draft refinement, preventing errors in unsupported claims or mismatched evidence. To mitigate errors within each agent’s workflow, the manager organizes feedback taskforces that maintain historical experience and implement dynamic checklists to guide self-correction before errors propagate to subsequent stages. Finally, the manager integrates the outputs of all taskforces to produce the final manuscript.
2 Related Work
@@
Automating the summarization of research papers has long attracted interest. Early work mainly used multi-document summarization techniques. In 2010, Hoang et al. introduced automatic related work summarization, extracting sentences from multiple articles to form an extractive summary tailored to the target paper Hoang and Kan (2010). Lu et al. constructed Multi-XScience, a large-scale dataset for generating related work sections based on a paper’s abstract and cited articles, serving as a benchmark for later studies Lu et al. (2020). Chen et al. proposed a relation-aware multi-document encoder and a relation graph to capture cross-document dependencies for abstractive related work generation Chen et al. (2021). Xiao et al. developed PRIMERA, a pre-trained encoder-decoder model with novel objectives for linking and aggregating information across documents Xiao et al. (2022). Liu et al. proposed aligning sentences by category and using sparse transformers to organize information from many references Liu et al. (2023).
@@
Recently, advances in LLMs have greatly improved the scope and quality of automated literature reviews, enabling end-to-end systems to generate literature review papers. Wang et al. introduced AutoSurvey, a framework with retrieval, outline generation, parallel subsection drafting, integration, and rigorous evaluation Wang et al. (2024b). Shao et al. developed STORM, which models the pre-writing stage by discovering diverse perspectives, simulating expert conversations, and curating information to create an outline used to generate a full Wikipedia article Shao et al. (2024). Liang et al. designed SurveyX, an end-to-end solution for automated survey generation, covering online literature search, organization, and survey writing Liang et al. (2025). While these methods represent impressive progress, little research has addressed the risk of errors in generated outputs. Self-correction algorithms in LLMs are not suitable for the workflow of literature review generation Madaan et al. (2023); Shinn et al. (2023); Ferraz et al. (2024). As systems become more complex, error accumulation becomes a significant challenge.
3 Methodology
@@ 3.1 Exploration Taskforce
In detail, given a user instruction $U$, the manager $A_{M}$ initiates the exploration taskforce. The manager constructs a tree with the user instruction as the root node, representing the overall direction for exploring the field. This tree is expanded by identifying sub-directions of interest. For the root node, $A_{M}$ first invokes the searching agent to retrieve literature based on the user instruction, resulting in a set of relevant papers. The searching agent $A_{S}$ operates as follows:
@@
Next, the manager assigns the outlining agent to determine sub-directions based on the titles and abstracts of the retrieved literature. The outlining agent $A_{O}$ operates as follows:
@@
For each outline node $O_{j}^{(d)}$ at depth $d$, the manager decides whether further decomposition is needed according to the following criterion:
@@ 3.2 Exploitation Taskforce
In detail, given the outline tree $\mathcal{T}$ produced as mentioned above, the manager $A_{M}$ initiates the exploitation taskforce. For each leaf node in the tree, $A_{M}$ identifies its parent outline node, denoted as $\{\mathcal{O}_{j}\}$, and the associated literature set $\{\mathcal{L}_{i}\}$. The exploitation taskforce then proceeds through iterative cycles. At the initial iteration $t=0$, the locating agent $A_{L}$ extracts relevant information based on the outline node $\{\mathcal{O}_{j}\}$ and the literature set $\{\mathcal{L}_{i}\}$, and operates as follows:
@@
where $\psi(\cdot,\cdot)$ denotes the similarity between consecutive iterations of the draft, measured by the ROUGE-1 score Lin (2004), $\epsilon$ is a similarity threshold that indicates no further significant improvements can be made, and $t_{\max}$ is the maximum number of allowed iterations.
The final content $M^{(t^{*})}_{i}$, where $t^{*}$ denotes the stopping iteration for parent node $i$, serves as the final draft. Finally, the manager $A_{M}$ aggregates the finalized drafts for all sections and delivers the completed literature review to the user.
3.3 Feedback Taskforce
The feedback taskforce is designed to execute tasks assigned by the manager, taking into account the risk of deviations at each step arising from unpredictable agent-environment interactions. To mitigate compounding errors, we propose an experience-based strategy.
In detail, for each agent $A_{W}$, the manager agent $A_{M}$ initiates the feedback taskforce. The agent first generates the primary results, then $A_{M}$ leverages historical experience to review these results, after which $A_{W}$ refines its outputs based on the feedback provided. At the initial step, $A_{W}$ produces the first output solely based on the context:
@@ 3.3 Feedback Taskforce
where $Y^{(0)}$ is the initial output, and $C$ is the context fetched by the manager agent.
[…]
@@
To give useful feedback, we propose an experience-aware mechanism. We record feedback and revision history in a database and select useful feedback for demonstration to provide feedback on the worker’s output at each iteration $x$.
@@
where useful feedback is defined as the feedback that has led to the fewest further revisions, where $\mathcal{R}_{\min}=\{r\in\mathcal{R}:\text{revision}(r)=r_{\min}\}$. Here, $\text{revision}(r)$ denotes the number of revisions after using feedback $r$. In $\mathcal{R}_{\min}$, we select the most useful one, i.e., with the highest $\text{usage}(r)$. Here, $\text{usage}(r)$ denotes the number of times the feedback $r$ is selected for demonstration.
Referring to $r^{*}$, the manager agent then provides feedback on the worker’s output:
@@
where $R^{(x)}$ denotes the feedback at iteration $x$.
For subsequent iterations ($x\geq 1$), the worker agent refines the output with the input and the feedback.
@@
The feedback taskforce continues until the output meets the manager’s requirements and no further feedback is provided, or until a maximum number of iterations $x_{\max}$ is reached. In the following, we detail how workers cooperate with the manager.
The searching agent formulates search queries based on user instructions, submits queries to search engines, and collects relevant literature. It filters and ranks the retrieved references by citation count, removes incomplete entries, and selects the top $Q_{k}$ unique references for further use. The outlining agent utilizes the gathered literature and user instructions to prompt an LLM for generating sub-outlines. The locating agent downloads and parses full-text papers, applies LLM-based previewing to select pertinent pages, and extracts concise text snippets for drafting. The drafting agent prompts the LLM to synthesize coherent text by incorporating selected references and prior drafts.
4 Experiments
4.1 Experimental Settings
We conduct experiments on two benchmark datasets, AutoSurvey Wang et al. (2024b) and SurveyEval Wang et al. (2025b), as well as a self-constructed benchmark dataset, TopSurvey. For hyperparameters, we set $d_{\max}$, $t_{\max}$, and $x_{\max}$ to 4. The hyperparameter $\theta$ is set to 500. We use GPT-4.1 for all agents. The evaluation comprises two categories of metrics. For citation quality, we adopt citation recall and citation precision as proposed by Wang et al. (2024b): recall measures whether cited passages fully support all statements, while precision measures the proportion of relevant citations that support their corresponding statements. For content quality, following Wang et al. (2024b), we use coverage, structure, and relevance, each rated by LLMs on a 5-point scale. Coverage assesses the extent to which the survey addresses all relevant aspects of the topic; structure evaluates logical organization and coherence; and relevance measures alignment with the specified research topic. We do not filter out fractional scores, such as 4.5.
4.2 Comparision Experiments